Let V and W be vector spaces over F, where V is n-dimensional. Let K ≤ V and R ≤ W

be finite-dimensional subspaces such that dim K + dim R = n. Prove that there exists a linear

transformation L : V −→ W such that ker L = K and Im L = R.
1
Expert's answer
2021-07-26T14:42:12-0400

Let "r=\dim R", then "\dim K=n-r".

Let "e_1, ..., e_r" be any basis for "R".

This basis can be extended to a basis "e_1, \dots, e_r, e_{r+1}, \dots, e_m" for "W", where "m=\dim W".

Let "f_{r+1}, ..., f_{n}" be any basis for "K".

This basis can be extended to a basis "f_1, \dots, f_{r}, f_{r+1}, \dots,f_n" for "V".

Define a linear transformation "L:V\to W" by the rule:

"L(a_1f_1+a_2f_2+\dots+a_nf_n)=a_{1}e_1+\dots +a_{r}e_{r}"


1) Show that "{\rm Im}\,L=R":

Since "e_1, ..., e_{r}" is a basis for "R",

"L(a_1 f_1+\dots+a_n f_n)=a_1e_1+\dots +a_{r}e_{r}\in R" and "{\rm Im}\,L\subset R"

For any "y\in R" if "y=y_1e_1+\dots + y_{r}e_{r}" then, evidently, "y=L(y_1f_1+\dots + y_{r}f_{r})\in {\rm Im}\,L".

Therefore, "{\rm Im}\,L=R".


2) Show that "{\rm Ker}\ L=K":

For any "x\in K" if "x=x_{r+1}f_{r+1}+\dots+x_nf_n" then "L(x)=0", i.e. "x\in{\rm Ker}\ L" and, hence, "K\subset {\rm Ker}\ L".

Conversely, if "L(x)=0" and "x=x_{1}f_{1}+\dots+x_nf_n" then "L(x)=x_{1}e_1+\dots +x_{r}e_{r}=0". Since Since "e_1, ..., e_{r}" is a basis for "R", they are linear independent, therefore "x_1=\dots=x_r=0".

This implies that "x=x_{r+1}f_{r+1}+\dots+x_nf_n\in K" and, hence, "{\rm Ker}\ L\subset K".

Therefore "{\rm Ker}\ L= K".